OpenStudy (anonymous):

Slove by the substittion method 4x5y=-4 7x+y=46

OpenStudy (anonymous):

4x5y = 20xy = 4 xy = 4/40 = 1/10 x = 1/10 / y = 1/10y

OpenStudy (anonymous):

You should be able to solve it now.

OpenStudy (anonymous):

pratu it is -4

OpenStudy (anonymous):

sorry, then its -1/10y

OpenStudy (anonymous):

y is in the denominator.

OpenStudy (saifoo.khan):

check your first equation @ asker. is it -ve or +ve?

OpenStudy (anonymous):

\[y = \frac{-1}{5x}\]

OpenStudy (anonymous):

i feel first equation should be 4x(+ or, -) 5y = -4

OpenStudy (anonymous):

plz cross check

OpenStudy (anonymous):

I also have 1/10

OpenStudy (anonymous):

-6 ,4 was the answer